Temperature and Quantum Number Dependence of Broadening Coefficients in the ν_{3} Band of SO_{2}

\underline{B. SUMPF*}, M. SCH NE, H.-D. KRONFELDT, Optisches Institut der Technischen Universit t Berlin, Sekr. PN 0-1, Hardenbergstr. 36, 10623 Berlin, Germany.

Systematic investigations concerning the dependence of the self-broadening coefficient γ_{SO_{2}−SO_{2}} on the quantum number J′′ and K′′_{a} are presented for 23 lines with 8 ≤ J′′ ≤ 33 and 1 ≤ K′′_{a} ≤ 16 at room temperature applying a diode laser spectrometer. We found that γ_{SO_{2}−SO_{2}} shows a typical decrease with increasing K′′_{a} whereas the J′′ dependence is not typical as in the case of the ν_{1} band. For the same lines collisions with air and hydrogen were studied. As in the case of self-broadening we found a similar behavior in the quantum number dependence. To complete our studies we measured the temperature dependence of the line broadening coefficients in the range between 250K and 360K with the perturbers air, helium, neon, argon, krypton, and xenon. For three lines with J′′ = 8 and K′′_{a} = 1, 2, 3 the noble gas broadening was studied whereas seven lines with 8 ≤ J′′ ≤ 30 and 1 ≤ K′′_{a} ≤ 9 were measured in the case of air broadening. The temperature coefficient n ranges between 0.56 and 0.89 for the noble gases and between 0.72 and 0.99 for air. For the noble gases one can suppose a weak decrease of n with rising atomic weight of the perturber although the experimental uncertainties are relatively large due to the restricted temperature range.

* B.S. acknowledges a grant from the Deutsche Forschungsgemeinschaft